JuliaGurl25:

How would the sum of cubes formula be used to factor x3y3 + 343? Explain the process. Do not write the factorization.

xXAikoXx:

The expression can be wirtten as, x^3y^3 + 343 = (xy)^3 +7^3 Therefore the given expression should be written as the cube's sum of xy and 7. A is = to x and y and b is also equal to the number 7. So, the expression is expressed as a sum for the cubes and it could be written as (xy)^3. The factors of the expression would approximately be (xy+7)(x^2y^2+49-7xy).

surjithayer:

\[a^3+b^3=\left( a+b \right)\left( a^2-ab+b^2 \right)\]